Describe what is meant by the fusion splicing of optical fibers. Discuss the advantages and drawbacks of this jointing technique. A multimode step index fiber with a core refractive index of 1.52 is fusion spliced. The splice exhibits an insertion loss of 0.8 dB. This insertion loss is found to be entirely due to the angular misalignment of the fiber core axes which is 7°. Determine the numerical aperture of the fiber.


1
Expert's answer
2021-04-15T20:41:17-0400



"\bigstar" Fusion splicing of optical fibers


"\bull" Splicing any fiber by making use of the fusion technique provides a permanent (long-lasting) contact between the two fibers. In the fusion splicing,


"\bull" the two fibers are thermally joined together. In this particular technique, an electrical instrument is necessarily used, that acts as an electric arc so as to form a thermal connection between the two.




"\bigstar" Diagram







Advantages splicing of optical fibers


  1. Fusion splicing offers a lower variable cost per fusion splice.
  2. It offers lower insertion loss and provides better performance, typical insertion loss is < 0.1 dB, therefore, it has very low impact on overall link performance.
  3. It provides better, more secure environment, so that there are minimum cable failures and weak signals.
  4. It provides a solution that is very compact and neat.
  5. It provides the lowest back reflection.





Disadvantages splicing of optical fibers

  1. The upfront cost of fusion splicing is very high mainly because of expensive fusion splicing device.
  2. Fusion splicing requires a constant power supply and some special tools.
  3. There are some situations where fusion splicing is not practical or we can say it is impossible, so as an alternative, we have to use mechanical splicing.
  4. Fusion splicing is a time consuming process and can’t be used for temporary connections.
  5. Fusion splicers need periodic maintenance which involves regular cleaning, electrode alignment and occasional replacement.




Numerical aperture 


"N_A= nsin(\alpha)"


"N_a=" numerical aperture


n= rifrective index


"\alpha" = angle


"N_a=" 1.52sin(7"^o" )

= 1.52"\times 0.12186934"


numerical aperture

"\\boxed{N_a\n\u200b\t\n = 0.1852413968}"